# Quiet Feature Learning in Algorithmic Tasks

### v2

@@ Abstract
We train Transformer-based language models on ten foundational algorithmic tasks and observe pronounced phase transitions in their loss curves that deviate from established power-law scaling trends. Over large ranges of compute, the validation loss barely improves, then abruptly decreases. Probing the models’ internal representations reveals that quiet features are learned prior to any decrease in task loss. These quiet features represent intermediate algorithmic computations that do not by themselves improve the output loss. Ablation experiments demonstrate that individual quiet features are causally necessary for task performance. Our results demonstrate that substantial representational progress can remain hidden beneath an apparently flat loss curve, challenging the prevailing use of cross-entropy as a proxy for learning and motivating richer diagnostics for monitoring model training.
11footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Understanding how and when large language models acquire new capabilities has become an important question in deep learning. While language models demonstrated remarkable performance across a broad range of tasks, the precise mechanisms driving their improvements remain unknown. Recent discussions of “emergent abilities" – where larger-scale models outperform baselines abruptly, even though smaller-scale counterparts exhibit little improvement – have led to debate over whether such phenomena are genuine or artifacts of measurement (Wei et al., 2022; Ganguli et al., 2022; Schaeffer et al., 2023).
Questions about emergent abilities are closely tied to the observation of scaling laws in model training (Kaplan et al., 2020; Ruan et al., 2024; Henighan et al., 2020; Dubey et al., 2024; OpenAI, 2023). These scaling laws typically show a smooth, power-law relationship between compute and model performance. However, most empirical demonstrations of these laws derive from heterogeneous data and tasks, leaving open the possibility that “averaging out" many distinct learning behaviors masks more abrupt transitions that occur for individual skills or subtasks.
@@
Figure 1: Model performance (validation loss) abruptly improves as we increase the model size, dataset size, and amount of compute (Training FLOPs) used for training. The input size for addition, multiplication, and activity selection is 16. For graph tasks, the input size is 11. For maximum subarray, the input size is 64, while for majority of majorities it is 32. The red dotted line indicates random performance.
@@ 2 Related Work
2.1 Scaling Laws
Hestness et al. (2017) observed that scaling dataset size and deep neural network model size led to a predictable decrease in generalization error for neural machine translation, language modeling, image classification and speech recognition. Kaplan et al. (2020) and Hoffmann et al. (2022) observed predictable relationships between training compute and language modeling loss. Henighan et al. (2020) extended this work for generative models across modalities: image, video, multimedia image-text and math. They demonstrated classification loss and error rates predictably decreased on downstream image classification tasks. Chen et al. (2021) studied language model performance on coding. They observed a predictable relationship between language modeling loss on a held out code corpus and model size.
2.2 Predicting LLM abilities
Ganguli et al. (2022) and Wei et al. (2022) demonstrated some large language models’ capabilities could not be predicted from capabilities of small language models. However, Schaeffer et al. (2023), OpenAI (2023), Ruan et al. (2024), and Dubey et al. (2024) provide evidence that this is due to choice of metrics and that large language models capabilities can be predicted from small language models.
2.3 Proposed explanations for scaling laws
Michaud et al. (2023) proposed that neural networks learn discrete skills called “quanta." They argue that there is a strict ordering, which they called Q sequence, in which quanta must be learned, and that the frequencies of these quanta follow a power law, leading to the power law relationship observed by Kaplan et al. (2020) and others. Hutter (2021) proposes that the relationship between the error rate and dataset size is guided by the distribution of features in the data. They show that a Zipfian distribution of features results in power law scaling.
2.4 Grokking
In grokking (Power et al., 2022; Nanda et al., 2023; Varma et al., 2023), a model trained for many epochs quickly memorizes the training set (thus achieving high training accuracy early) but only later learns a generalizing solution, causing a sudden jump in test accuracy. Our scaling law results are related to grokking, but occur in the single epoch setting. Unlike grokking, models trained in the single-epoch setting do not exhibit a transition from memorization to generalization.
Figure 2: Models exhibit similar abrupt improvement in performance during a single training run. Plots show compute-optimal training runs for the smallest compute budget where test accuracy is 100%. The red dotted line indicates random performance.
2.5 Progress Measures
@@
2.6 Phase Transitions
Phase transitions were previously observed for a limited number of algorithmic tasks. Olsson et al. (2022), Garg et al. (2022), and Edelman et al. (2024) find phase transitions for in-context learning during individual training runs. Barak et al. (2022) observed phase transitions in parity. Lee et al. (2024) measure relationships between test accuracy and number of examples (over a fixed model size), with observed phase transitions potentially being explained by the metric artifacts of Schaeffer et al. (2023).
@@ 3.1 Task Formulation
All tasks are formulated as sequence prediction problems. The input to the problem is serialized, and an autoregressive model is trained to predict the solution. All tasks use a standard cross-entropy loss, with the loss masked on the input tokens. We describe how we formulate three of the tasks below. For other tasks, please see Appendix A.
3.1.1 Binary Addition
@@
3.1.2 Breadth First Search
@@ 3.1.3 Maximum Subarray
3.2 Experimental Methodology
Figure 3: Models learn quiet features before the phase transition. The loss is averaged over the first third of token positions (Beginning), second third (Middle), and last third (End). The red vertical line indicates the task success threshold, which is the smallest compute budget at which the task loss starts to decrease (see Appendix Figure 6). Horizontal dotted lines represent random baselines.
3.2.1 Model Training
Each task is trained independently with the Transformer++ architecture. Transformer++ (Gu and Dao, 2024) is a decoder-only transformer model with enhancements detailed in Appendix Table 4, based on modifications in Llama and PaLM (Touvron et al., 2023; Chowdhery et al., 2023). This architecture is chosen because it has improved performance in scaling law experiments compared to other transformer variants (Gu and Dao, 2024). Models are trained with the AdamW optimizer (Loshchilov and Hutter, 2017) with linear warmup followed with cosine learning rate annealing as prescribed by Hoffmann et al. (2022).
3.2.2 Estimating Scaling Laws
The scaling law experiments aim to estimate the best performance achievable on a task given a compute budget. Separate scaling laws are estimated for each task and input size. Each model is trained up to a pre-specified compute budget, which ranged from $10^{9}-10^{15}$ FLOPs. [^1] For multiplication, the maximum budget was increased to $10^{16}$, since this was the minimum budget needed to train the task to 100% accuracy. For each budget, we conduct a grid search across model sizes, batch sizes, and learning rates (see Appendix Table 3 for details about the hyperparameter search). Following the procedure from Chinchilla (Hoffmann et al., 2022), the period of the learning rate scheduler is set to the number of training steps.
The number of training runs per task varies from 1316 to 3565, and the total number of training runs is 18544. All models are trained for at most a single epoch; each algorithmic task has a sufficient number of unique examples to avoid repetition even with the highest compute budgets. The number of training examples is determined based on training compute budget and model size, with all configurations evaluated using randomly generated validation and test sets with 1000 examples each. We choose the configuration with minimum validation loss for each training compute and designate it as the “compute-optimal validation loss."
@@ 3.3 Scaling Law Results
Next, we investigate the effects of varying the dataset size. In Appendix Figure 6, we fix the model size (selecting the model size corresponding to the smallest training compute budget that achieves 100% test accuracy) and increase the dataset size. We continue to observe phase transitions even when the model size is fixed. In this setting, additional graph tasks exhibit distinct phase transitions.
@@
Phase transitions in compute-optimal validation losses occur across different task sizes (see Appendix Figure 7). For addition, phase transitions are observed across task sizes and similarly within individual training runs (Appendix Figure 9). As the input size increases, the Pareto frontier shifts to the right but maintains the same shape. However, for maximum subarray, the phase transition only appears at task sizes greater than 16.
Figure 5: Models learn different set of features (loud features) at or after the phase transition. The loss is averaged over the first third of token positions (Beginning), second third (Middle), and last third (End). The red vertical line indicates the task success threshold, which is the smallest compute budget at which the task loss starts to decrease (see Appendix Figure 6). Cross-entropy loss is used for training probes for first_operand, adjacency_list. The probing loss is mean squared error for start_time and max_ending_here, since these are continuous features.
@@ 4 Feature Learning before Phase Transitions
4.1 Feature Probing Methodology
For each algorithm-specific feature, we train separate linear probes across each token position and each layer. Probes are trained on the residual streams after each layer (see Appendix B.3). Each probe is trained with 10,000 examples which had been held-out from the original model training set.
@@ 4.1 Feature Probing Methodology
4.2 Intermediate Task Features
@@
Table 1: Average difference in test accuracy after ablating a quiet feature compared to ablating a random direction (random ablation). Ablating quiet features degrades test accuracy more than random ablation. For random ablation, we estimate test accuracy over 32 trials. indicates $p<0.001$ using bootstrapping. For complete accuracy / loss values see Appendix Table 6
4.3 Feature Probing Results
The model learns algorithmic features before, during and after the phase transition. We call features learned prior to the phase transition quiet features, as they occur during the slow phase where loss is stagnant or slowly decreasing. Features learned in the fast phase (during and after the phase transition) are loud features as the task loss decreases rapidly. Figure 3 shows the trajectory of the probing loss for quiet features. For addition, multiplication, and maximum subarray, the model learns features for early token positions prior to the phase transition. However, for breadth first search, later token positions are learned first.
@@ 4.3 Feature Probing Results
Figure 5 shows models learn loud features in the fast phase (during and after the phase transition).
A surprising finding is the U-shaped feature learning curves in Figure 3, indicating that the probing loss increases for many quiet features after the phase transition. This may indicate that the models are learning alternative representations in the highest compute budget regimes, though the probing loss remains below the random baseline.
4.4 Are quiet features causal?
@@
Ablation results are shown in Table 1. Quiet feature ablations are compared to ablating a random direction. When we ablate quiet features, we observe test accuracy generally degrades more than ablating a random direction, indicating a causal role for quiet features. Similar results are seen for test loss, as shown in Appendix Table 6. However, for first_operand at input size 16, we do not see any significant change after ablating the feature compared to random in addition or multiplication. At larger input sizes, ablating first_operand for addition leads to significant test loss degradation compared to random ablation.
@@ 5 Discussion
These findings have practical and conceptual implications. For practitioners, they highlight the risk of judging model capabilities based solely on loss curves. Probe or circuit-based diagnostics could provide earlier warnings that a model is nearing a capability threshold. Conceptually, they raise questions about whether similar quiet phases exist in more complex natural-language settings. They also underscore the need for theoretical frameworks that explain why models accumulate latent subroutines before they begin to pay off in observable metrics.
@@ Appendix A Task formulation
Given a connected undirected graph $G$ with $n$ vertices and a set of weighted edges $E$, we predict the set of edges forming the minimum spanning tree (MST).
@@ B.1 Generating Samples
For graph-based tasks (breadth first search, depth first search, shortest path, minimum spanning tree and topological sorting), we uniformly sample non-isomorphic undirected connected graphs, using the graph dataset of McKay [2025], and randomly permute the vertex labels. For topological sorting, edge directions are determined by randomly sampling a vertex ordering.
@@
B.2 Estimating Scaling Laws (Additional Details)
@@
B.3 Training Feature Probes
B.3.1 Transformer Architecture and Residual Stream
@@
B.3.2 Probe Training Procedure
@@
For continuous features, max_ending_here & start_time, we train a linear regressor:
@@ B.3.4 Probe Selection
Figure 7: Models trained for addition exhibit phase transition for all task lengths. However, models trained for maximum subarray do not exhibit phase transition for smaller task lengths
@@
Figure 9: Models exhibit phase transitions for increasing task lengths. Compute-optimal training run is selected for the earliest training FLOPs in the same fashion as Figure 2

### v3

@@ Abstract
We train Transformer-based language models on ten foundational algorithmic tasks and observe pronounced phase transitions in their loss curves that deviate from established power-law scaling trends. Over large ranges of compute, the validation loss barely improves, then abruptly decreases. Probing the models’ internal representations reveals that quiet features are learned prior to any decrease in task loss. These quiet features represent intermediate algorithmic computations that do not by themselves improve the output loss. Ablation experiments demonstrate that individual quiet features are causally necessary for task performance. Our results demonstrate that substantial representational progress can remain hidden beneath an apparently flat loss curve, challenging the prevailing use of cross‑entropy as a proxy for learning and motivating richer diagnostics for monitoring model training.
Codehttps://github.com/prudhvirajn/quiet-feature-learning-in-algorithmic-tasks
@@ 1 Introduction
Understanding how and when large language models acquire new capabilities has become an important question in deep learning. While language models demonstrated remarkable performance across a broad range of tasks, the precise mechanisms driving their improvements remain unknown. Recent discussions of “emergent abilities” – where larger-scale models outperform baselines abruptly, even though smaller-scale counterparts exhibit little improvement – have led to debate over whether such phenomena are genuine or artifacts of measurement (Wei et al. 2022; Ganguli et al. 2022; Schaeffer, Miranda, and Koyejo 2023).
Questions about emergent abilities are closely tied to the observation of scaling laws in model training (Kaplan et al. 2020; Ruan, Maddison, and Hashimoto 2024; Henighan et al. 2020; Dubey et al. 2024; OpenAI 2023). These scaling laws typically show a smooth, power-law relationship between compute and model performance. However, most empirical demonstrations of these laws derive from heterogeneous data and tasks, leaving open the possibility that “averaging out” many distinct learning behaviors masks more abrupt transitions that occur for individual skills or subtasks.
@@
Figure 1: Model performance (validation loss) abruptly improves as we increase the model size, dataset size, and amount of compute (Training FLOPs) used for training. The input size for addition, multiplication and activity selection is 16. For graph tasks, the input size is 11. For maximum subarray, the input size is 64 while for majority of majorities it is 32. The red dotted line indicates random performance.
@@ 2 Related Work
Scaling Laws
Hestness et al. (2017) observed scaling dataset size and deep neural network model size led to a predictable decrease in generalization error for neural machine translation, language modeling, image classification and speech recognition. Kaplan et al. (2020) and Hoffmann et al. (2022) observed predictable relationships between training compute and language modeling loss. Henighan et al. (2020) extended this work for generative models across modalities: image, video, multimedia image-text and math. They demonstrated classification loss and error rates predictably decreased on downstream image classification tasks. Chen et al. (2021) studied language model performance on coding. They observed a predictable relationship between language modeling loss on a held out code corpus and model size.
Predicting LLM abilities
Ganguli et al. (2022) and Wei et al. (2022) demonstrated some large language models’ capabilities could not be predicted from capabilities of small language models. However, Schaeffer, Miranda, and Koyejo (2023), OpenAI (2023), Ruan, Maddison, and Hashimoto (2024), and Dubey et al. (2024) provide evidence that this is due to choice of metrics and that large language models capabilities can be predicted from small language models.
Proposed explanations for scaling laws
Michaud et al. (2023) proposed neural networks learn discrete skills called “quanta.” They argue that there is a strict ordering, which they called Q sequence, in which quanta must be learned, and that the frequencies of these quanta follow a power law, leading to the power law relationship observed by Kaplan et al. (2020) and others. Hutter (2021) propose that the relationship between the error rate and dataset size is guided by the distribution of features in the data. They show a Zipfian distribution of features results in power law scaling.
Grokking
In grokking (Power et al. 2022; Nanda et al. 2023; Varma et al. 2023), a model trained for many epochs quickly memorizes the training set (thus achieving high training accuracy early) but only later learns a generalizing solution, causing a sudden jump in test accuracy. Our scaling law results are related to grokking, but occur in the single epoch setting. Unlike grokking, models trained in the single-epoch setting do not exhibit a transition from memorization to generalization.
Figure 2: Model exhibit similar abrupt improvement in performance during a single training run. Plots show compute-optimal training runs for the smallest compute budget where test accuracy is 100%. The red dotted line indicates random performance.
Progress Measures
@@
Phase Transitions
Phase transitions were previously observed for a limited number of algorithmic tasks. Olsson et al. (2022), Garg et al. (2022), and Edelman et al. (2024) find phase transitions for in-context learning during individual training runs. Barak et al. (2022) observed phase transitions in parity. Lee et al. (2024) measures relationships between test accuracy and number of examples (over a fixed model size), with observed phase transitions potentially being explained by the metric artifacts of Schaeffer, Miranda, and Koyejo (2023).
@@ 3 Scaling Laws for Algorithmic Tasks
All tasks are formulated as sequence prediction problems. The input to the problem is serialized, and an autoregressive model is trained to predict the solution. All tasks use a standard cross-entropy loss, with the loss masked on the input tokens. We describe how we formulate three of the tasks below. For other tasks, please see Technical Appendix A.
Binary Addition
@@
Breadth First Search
@@
Experimental Methodology
Figure 3: Models learn quiet features before the phase transition. The loss is averaged over the first third of token positions (Beginning), second third (Middle), and last third (End). The red vertical line indicates the task success threshold, which is the smallest compute budget at which the task loss starts to decrease (see Technical Appendix Figure 6). Horizontal dotted lines represent random baselines.
Model Training
Each task is trained independently with the Transformer++ architecture. Transformer++ is a decoder-only transformer model with enhancements detailed in Technical Appendix Table 4, based on modifications in Llama and PALM (Gu and Dao 2024). This architecture is chosen because it has improved performance in scaling law experiments compared to other transformer variants (Gu and Dao 2024).
Models are trained with the AdamW optimizer (Loshchilov and Hutter 2019) with linear warmup followed with cosine learning rate annealing as prescribed by Hoffmann et al. (2022).
Estimating Scaling Laws
The scaling law experiments aim to estimate the best performance achievable on a task given a compute budget. Separate scaling laws are estimated for each task and input size. Each model is trained up to a pre-specified compute budget, which ranged from $10^{9}-10^{15}$ FLOPs. [^1] For multiplication, the maximum budget was increased to $10^{16}$, since this was the minimum budget needed to train the task to 100% accuracy. For each budget, we conduct a grid search across model sizes, batch sizes, and learning rates (see Technical Appendix Table 3 for details about the hyperparameter search). Following the procedure from Chinchilla (Hoffmann et al. 2022), the period of the learning rate scheduler is set to the number of training steps.
The number of training runs per task varies from 1316 to 3565, and the total number of training runs is 18544. All models are trained for at most a single epoch; each algorithmic task has a sufficient number of unique examples to avoid repetition even with the highest compute budgets. The number of training examples is determined based on training compute budget and model size, with all configurations evaluated using randomly generated validation and test sets with 1000 examples each. We choose the configuration with minimum validation loss for each training compute and designate it as the “compute-optimal validation loss.”
@@
Next, we investigate the effects of varying the dataset size. In Technical Appendix Figure 6, we fix the model size (selecting the model size corresponding to the smallest training compute budget that achieves 100% test accuracy) and increase the dataset size. We continue to observe phase transitions even when the model size is fixed. In this setting, additional graph tasks exhibit distinct phase transitions.
@@
Phase transitions in compute-optimal validation losses occur across different task sizes (see Technical Appendix Figure 8). For addition, phase transitions are observed across task sizes and similarly within individual training runs (Technical Appendix Figure 9). As the input size increases, the Pareto frontier shifts to the right but maintains the same shape. However, for maximum subarray, the phase transition only appears at task sizes greater than 16.
Figure 5: Models learn different set of features (loud features) at or after the phase transition. The loss is averaged over the first third of token positions (Beginning), second third (Middle), and last third (End). The red vertical line indicates the task success threshold, which is the smallest compute budget at which the task loss starts to decrease (see Technical Appendix Figure 6). Cross-entropy loss is used for training probes for first_operand, adjacency_list. The probing loss is mean squared error for start_time and max_ending_here, since these are continuous features.
@@ 4 Feature Learning before Phase Transitions
Feature Probing Methodology
For each algorithm-specific feature, we train separate linear probes across each token position and each layer. Probes are trained on the residual streams after each layer (See Technical Appendix B). Each probe is trained with 10,000 examples which had been held-out from the original model training set.
@@
Intermediate Task Features
@@
Table 1: Average difference in test accuracy after ablating a quiet feature compared to ablating a random direction (random ablation). Ablating quiet features degrades test accuracy more than random ablation. For random ablation, we estimate test accuracy over 32 trials. indicates $p<0.001$ using bootstrapping. For complete accuracy / loss values see Technical Appendix Table 6
Feature Probing Results
The model learns algorithmic features before, during and after the phase transition. We call features learned prior to the phase transition quiet features, as they occur during the slow phase where loss is stagnant or slowly decreasing. Features learned in the fast phase (during and after the phase transition) are loud features as the task loss decreases rapidly. Figure 3 shows the trajectory of the probing loss for quiet features. For addition, multiplication, and maximum subarray, the model learns features for early token positions prior to the phase transition. However, for breadth-first search, later token positions are learned first.
@@
Figure 5 shows models learns loud features in the fast phase (during and after the phase transition).
A surprising finding is the U-shape of feature learning curves in Figure 3, indicating that the probing loss increases for many quiet features after the phase transition. This may indicate that the models are learning alternative representations in the highest compute budget regimes, though the probing loss remains below the random baseline.
Are quiet features causal?
@@
Ablation results are shown in Table 1. Quiet feature ablations are compared to ablating a random direction. When we ablate quiet features, we observe test accuracy generally degrades more than ablating a random direction, indicating a causal role for quiet features. Similar results are seen for test loss, as shown in Technical Appendix Table 6. However, for first_operand at input size 16, we do not see any significant change after ablating the feature compared to random in addition or multiplication. At larger input sizes, ablating first_operand for addition leads to significant test loss degradation compared to random ablation.
@@ 5 Discussion
These findings have practical and conceptual implications. For practitioners, they highlight the risk of judging model capabilities based solely on loss curves. Probe or circuit-based diagnostics could provide earlier warnings that a model is nearing a capability threshold. Conceptually, they raise questions about whether similar quiet phases exist in more complex natural‑language settings. They also underscore the need for theoretical frameworks that explain why models accumulate latent subroutines before they begin to pay off in observable metrics.
@@ Appendix A Task formulation
Given a connected undirected graph $G$ with $n$ vertices and a set of weighted edges $E$, we predict the set of edges forming the minimum spanning tree (MST)..
@@ Appendix B Experimental Methodology
For graph-based tasks (breadth first search, depth first search, shortest path, minimum spanning tree and topological sorting), we uniformly sample non-isomorphic undirected connected graphs, using the graph dataset of McKay (2025), and randomly permute the vertex labels. For topological sorting, edge directions are determined by randomly sampling a vertex ordering.
@@
Estimating Scaling Laws (Additional Details)
@@
Training Feature Probes
Transformer Architecture and Residual Stream
@@
Probe Training Procedure
@@
For continuous features, max_ending_here & retrieve_start_times, we train a linear regressor:
@@
Figure 7: Validation loss for best run from grid search vs training FLOPs for addition & maximum subarray across different input sizes
@@
Figure 9: Validation loss vs training steps for a single training run for addition across input sizes. Compute-optimal training run is selected for the earliest training FLOPs in the same fashion Figure 2